The following information applies to the products of Campbell Company.

Product A Product B
Selling price per unit $ 12 $ 18
Variable cost per unit 8 12


Identify the product that should be produced or sold under each of the following constraints. Consider each constraint separately.

Required

  1. One unit of Product A requires 2 hours of labor to produce, and one unit of Product B requires 5 hours of labor to produce. Due to labor constraints, demand is higher than the company’s capacity to make both products.

  2. The products are sold to the public in retail stores. The company has limited floor space and cannot stock as many products as it would like. Display space is available for only one of the two products. Expected sales of Product A and Product B are 13,000 units and 11,000 units, respectively.

  3. The maximum number of machine hours available is 42,000. Product A uses 2 machine hours, and Product B uses 6 machine hours. The company can sell all the products it produces.
Answer #1

a.

Contribution margin per direct labour hour:

Product A = ($12 - $8) / 2 = $2

Product B = ($18 - $12) / 5 = $1.20

Product A has more Contribution margin per direct labor hour. Product A should be produced.

b.

Total contribution margin:

Product A = $4 X 13,000 = $52,000

Product B = $6 X 11,000 = $66,000

Product B has highest Total contribution margin. Product B should be produced.

c.

Contribution margin per direct labour hour:

Product A = ($12 - $8) / 2 = $2

Product B = ($18 - $12) / 6 = $1

Product A has more Contribution margin per direct labor hour. Product A should be produced.